OpenStudy (anonymous):

In 4 years, mary will be twice as old as she was last year. How old is mary now?

Parth (parthkohli):

If mary is x years old, then she was x - 1 last year. Agree?

OpenStudy (anonymous):

Agree

Parth (parthkohli):

So, x + 4 = 2(x - 1) Solve!

OpenStudy (wasiqss):

wa wa parth dimag chal raha hai :D

OpenStudy (wasiqss):

@baita

Parth (parthkohli):

Lol was thank you

OpenStudy (apoorvk):

Let Mary's present age be 'x' So her age an year ago = x-1 after four years, her age = x+4 So what would the equation for the relation between the two ages be?

Parth (parthkohli):

twice as old.......as she was last year

OpenStudy (apoorvk):

Solve that^

OpenStudy (anonymous):

6 years. Thanks @ParthKohli and @apoorvk

Parth (parthkohli):

TEEHEE!

OpenStudy (apoorvk):

Righty-right!